# Symmetry-breaking induced surface magnetization in non-magnetic RuO2

### v1

@@ III.2.3 The significance of surface magnetism in RuO2 (110) surface
Surface magnetization can also exhibit different characteristics when probed by surface- and spin-sensitive microscopy techniques such as spin-polarized scanning tunneling microscopy (SP-STM) [55]. Since SP-STM is applicable to electrically conductive materials, such as magnetic metals or doped semiconductors, RuO2’s metallic nature makes it a suitable candidate for the technique. General observation from STM images shown in Fig. 5 is the high intensity at bright spots corresponding to the O-br–derived electronic states due to their positions lying closest to the tip in STM measurement. This has been theoretically simulated and experimentally observed [56, 57]. Notably, although Ru-5f exhibits a high DOS for both spin channels below the Fermi level, the signal is predominantly dominated by bridging oxygen atoms rather than Ru-5f (cf. Fig. 5(a),(c) and Fig. S6(a),(b),(e),(f)). This occurs because the Ru-5f atoms are located one atomic layer further from the STM tip compared to the bridging oxygen atoms, resulting in an exponential decrease in tunneling current with distance. Furthermore, the occupied states of Ru-5f originate from in-plane Ru $4d$ orbitals (Fig. S6(i)), which do not extend significantly out of plane. This interpretation is further supported by spin density visualizations shown in Fig. S6(j),(l), revealing that, under negative bias voltage, the spin density of Ru-6f and O-br covers a greater spatial extent than that of the Ru-5f.
@@
Figure 5: Simulated STM images of (110) surface. (a)–(d) Images at bias voltages of $-0.5$ V (left column) and $+0.5$ V (right column); the unit cell of the surface highlighted by a dashed green rectangle; Ru-6f, Ru-5f and oxygen atoms represented by red, blue, and bright gray balls, respectively. (e)–(l) Bias- and spin-dependent line profiles indicated as lines 1, 2, 3, and 4 in (c), left and right column figures corresponding to $-0.5$ and $+0.5$ V bias voltages, respectively. (m) Site-projected DOS of surface atoms explaining the difference in spin-dependent images. Atomic labels denoted as in Fig. 2(b).
@@
More importantly, our simulated STM images shown in Fig. 5 reveal clear differences in the spin-dependent images. For instance, under negative bias voltage (i.e., probing occupied states), we can see high protrusions at Ru-6f and O-br positions forming horizontal lines in the spin-up images. This is due to the occupied states being dominated by the majority spin (spin-up) channel mostly derived from Ru-6f and O-br as identified by the atomic (site-) projected density of states for surface atoms (Fig. 5(m)) and the charge distribution from partial charge density calculations (Fig. S6(j)). In contrast, when probing the unoccupied states under positive bias voltage, higher protrusions forming horizontal lines in the obtained images were derived from the spin-down channel. The spin-dependent feature of the STM images can be clearly corroborated further by height–profile of the signals scanned over four selected lines indicated in Fig. 5(c). Figures in the left column corresponding to the negative bias voltage images show the dominant contribution of the spin-up density from Ru-6f and O-br positions. In contrast, the figures on the right column, which represent positive bias voltage scans, show the significant effect of the spin-down density from Ru-6f and O-br. This aligns well with the dominant contribution of majority and minority spin density of states in a spin-polarized system.
@@
An intriguing feature emerges when transitioning from negative to positive bias: higher positive biases enhance the visibility of the Ru-5f topography. At +0.5 V bias, the position of Ru-5f becomes discernible in the spin-up images; and at +1.0 V bias, bright vertical columns along [$01\overline{1}$ 0] appearing periodically along the [001] crystallographic direction are obvious(Fig. S6(c)). This effect arises from the dominant contribution of the Ru-5f spin-up channel over the spin-down channel in unoccupied states as demonstrated in Fig. 5(e). As the bias voltage increases, the position of Ru-5f atoms become more clearly visible due to the higher contribution from the out-of-plane components of its $4d$ orbitals as shown in Fig. S6(i), exhibiting a phenomenon known as contrast reversal. This behavior has previously been observed in studies of rutile transition metal oxide (110) surfaces and helps distinguish stoichiometric surfaces from other structures under varying experimental conditions, further validating our STM image simulations [57].
@@ IV Conclusions
In conclusion, we have shown by first-principles calculations that surface magnetization can be developed from a metallic non-magnetic bulk material. The magnetization is significantly large at the outermost layers of the slab structure and decays quickly as it gets deeper into the bulk region. Symmetry lowering as a result of surface termination induces not only structural relaxation but also significant electronic reorganization at the surface layer, resulting in sizable moments on the atoms forming the surface. This spontaneous surface magnetism proved to exhibit unique properties, including spin-split surface bands near the Fermi level and distinct spin-resolved spectroscopic features. These findings are expected to provide valuable insights into the interpretation of experimental observations in spectroscopy and spin-dependent transport phenomena, highlighting the pivotal role of surface magnetism in such systems.
@@ V Acknowledgement
This work was supported by the NSF through the UD-CHARM University of Delaware Materials Research Science and Engineering Center (MRSEC) Grant No. DMR-2011824. A.J. acknowledges support from the U.S. Department of Energy (Contract No. DE-SC0014388). We also acknowledge the use of Stampede3 at TACC through allocation PHY240154 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation Grants No. 2138259, 2138286, 2138307, 2137603, and 2138296, and the DARWIN computing system at the University of Delaware, which is supported by the NSF Grant No. 1919839.

### v2

@@ III.2.3 The significance of surface magnetism in RuO2(110)
Surface magnetization can also exhibit different characteristics when probed by surface- and spin-sensitive microscopy techniques such as spin-polarized scanning tunneling microscopy (SP-STM) [59]. Since SP-STM is applicable to electrically conductive materials, such as magnetic metals or doped semiconductors, RuO2’s metallic nature makes it a suitable candidate for the technique. General observation from STM images shown in Fig. 5 is the high intensity at bright spots corresponding to the O-br–derived electronic states due to their positions lying closest to the tip in STM measurement. This has been theoretically simulated and experimentally observed [60, 61]. Notably, although Ru-5f exhibits a high DOS for both spin channels below the Fermi level, the signal is predominantly dominated by bridging oxygen atoms rather than Ru-5f [(cf. Fig. 5(a),(c) and Fig. S6(a),(b),(e),(f)]. This occurs because the Ru-5f atoms are located one atomic layer further from the STM tip compared to the bridging oxygen atoms, resulting in an exponential decrease in tunneling current with distance. Furthermore, the occupied states of Ru-5f originate from in-plane Ru $4d$ orbitals (Fig. S6(i)), which do not extend significantly out of plane. This interpretation is further supported by spin density visualizations shown in Fig. S6(j),(l), revealing that, under negative bias voltage, the spin density of Ru-6f and O-br covers a greater spatial extent than that of the Ru-5f.
@@
Figure 5: Simulated STM images of (110) surface. (a)–(d) Images at bias voltages of $-0.5$ V (left column) and $+0.5$ V (right column); the unit cell of the surface highlighted by a dashed green rectangle; Ru-6f, Ru-5f and oxygen atoms represented by red, blue, and bright gray balls, respectively. (e)–(l) Bias- and spin-dependent line profiles indicated as lines 1, 2, 3, and 4 in (c), left and right column figures corresponding to $-0.5$ V and $+0.5$ V bias voltages, respectively. (m) Site-projected DOS of surface atoms explaining the difference in spin-dependent images. Atomic labels denoted as in Fig. 2(b).
@@
More importantly, our simulated STM images shown in Fig. 5 reveal clear differences in the spin-dependent images. For instance, under negative bias voltage (i.e., probing occupied states), we can see high protrusions at Ru-6f and O-br positions forming horizontal lines in the spin-up images. This is due to the occupied states being dominated by the majority spin (spin-up) channel mostly derived from Ru-6f and O-br as identified by the atomic (site-) projected density of states for surface atoms (Fig. 5(m)) and the charge distribution from partial charge density calculations (Fig. S6(j)). In contrast, when probing the unoccupied states under positive bias voltages, higher protrusions forming horizontal lines in the obtained images were derived from the spin-down channel.
The spin-dependent feature of the STM images can be clearly corroborated further by the height–profiles of the signals scanned over four selected lines indicated in Fig. 5(c). The profile lines displayed in the left-column figures [Fig. 5(e),(g),(i),(k)] corresponding to the negative bias voltage images exhibit periodic patterns consistent with the dominant contribution of the spin-up density from Ru-6f and O-br positions. In contrast, the right-column figures [Fig. 5(f),(h),(j),(l)], representing positive bias voltage scans, show the significant effect of the spin-down density from the same atomic positions. This aligns well with the expected contribution of the majority and minority spin density of states in a spin-polarized system.
@@
An intriguing feature emerges when transitioning from negative to positive bias: higher positive biases enhance the visibility of the Ru-5f topography. At +0.5 V bias, the position of Ru-5f becomes discernible in the spin-up images; and at +1.0 V bias, bright vertical columns along [$01\overline{1}$ 0] appearing periodically along the [001] crystallographic direction are obvious (Fig. S6(c)). This effect arises from the dominant contribution of the Ru-5f spin-up channel over the spin-down channel in its unoccupied states as demonstrated in Fig. 5(m) and Fig. S6(i). As the bias voltage increases, the position of Ru-5f atoms become more clearly visible due to the higher contribution from the out-of-plane components of its $4d$ orbitals as shown in Fig. S6(i), exhibiting a phenomenon known as contrast reversal. This behavior has previously been observed in studies of rutile transition-metal oxide (110) surfaces and helps distinguish stoichiometric surfaces from other structures under varying experimental conditions, further validating our STM image simulations [61].
@@ IV Conclusions
In summary, we have shown by first-principles calculations that surface magnetization emerges from a metallic non-magnetic bulk material. The magnetization is significantly large at the outermost layers of the slab structure and decays quickly as it gets deeper into the bulk region. Symmetry lowering as a result of surface termination induces not only structural relaxation but also significant electronic reorganization at the surface layer, resulting in sizable moments on the atoms forming the surface. This spontaneous surface magnetism proved to exhibit unique properties, including spin-split surface bands near the Fermi level and distinct spin-resolved spectroscopic features. These findings are expected to provide valuable insights into the interpretation of experimental observations in spectroscopy and spin-dependent transport phenomena, highlighting the pivotal role of surface magnetism in such systems.
@@ V Acknowledgement
We thank D. Wickramaratne, J. Xiao, S. Bhatt, and D. Plouff for fruitful discussions. This work was supported by the NSF through the UD-CHARM University of Delaware Materials Research Science and Engineering Center (MRSEC) Grant No. DMR-2011824. A.J. acknowledges support from the U.S. Department of Energy (Contract No. DE-SC0014388). We also acknowledge the use of Stampede3 at TACC through allocation PHY240154 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation Grants No. 2138259, 2138286, 2138307, 2137603, and 2138296, and the DARWIN computing system at the University of Delaware, which is supported by the NSF Grant No. 1919839.